OpenStudy (anonymous):

using substitution solve for the equation y=1/2x+3 y-5=x which one do i substitute?

OpenStudy (saifoo.khan):

y=1/2x+3 -1 y-5=x - 2 \ Solve 2. y = 5 + x Therefore, y=1/2x+3 5+x = 1/2x =3

OpenStudy (anonymous):

so there is no value for x or y in this problem?'

OpenStudy (saifoo.khan):

There is, solve this further; 5+x = 1/2x =3 and you will get the value for x. then use that value to find the value for y.

OpenStudy (anonymous):

what do i do next in order to solve this?

OpenStudy (saifoo.khan):

make x the subject.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

just subtract one x and move it to the other side right?

OpenStudy (zepp):

You can do this http://puu.sh/qKQg

OpenStudy (saifoo.khan):

\[5+x = \frac12x =3\]\[x - \frac12 x = -5 + 3\]

OpenStudy (anonymous):

okay then what

OpenStudy (saifoo.khan):

\[\frac{2x-1x}{2} = -2\]\[x = -4\]

OpenStudy (anonymous):

where did that come from??

OpenStudy (saifoo.khan):

i just solved with you step by step.

OpenStudy (anonymous):

oh ok so what would the value of y be then??

OpenStudy (zepp):

Then this http://puu.sh/qKS7

OpenStudy (saifoo.khan):

@zepp , are you using mouse to write that? :O

OpenStudy (saifoo.khan):

now use this value and insert it into any another equation of two given/ let's say second one, y-5=x Y - 5 = -4 y = -4 + 5 y = 1

OpenStudy (anonymous):

oh okay

OpenStudy (zepp):

I used a pen tablet :P

OpenStudy (saifoo.khan):

i want one as well. :/